if you live in a tent whose inside volume is 203 m3 and the temperature in a cold morning is 15.o degree C but becomes 20.0 degees C in the afternonn due to the sun's warming, how many cubic meter of air was forced out of the tent due to the sun's warming . Assume that the pressure in the tent remains constant.
1
Expert's answer
2012-03-29T11:22:22-0400
P1V2=nRT1
P2V2=nRT2
R=const
n=P1*V1/T1
and P2*V2/T2
P1*V1/T1=P2*V2/T2
if
P=const
V1/T1=V2/T2
V2=V1*T2/T1
V2=203 m3 *
(293K/288K)=206,5
V2-V1=206,5-203=3,5 L